OpenStudy (anonymous):

If it takes 44.97 mL of a 0.87 acid to neutralize 37 mL of a base, what was is the concentration of the base?

OpenStudy (anonymous):

This is jsut law of equivalents.N1V1=N2V2 44.97*0.87=37*x x=44.97*0.87/37=1.057 normal